OpenStudy (jemson):

(x+1/x)^2 + (x^2 +1/x^2)^2+………(x^27 + 1/x^27)^2 = ?

OpenStudy (jemson):

If G is the geometric mean of X & Y, then (1/(g^2-x^2)) + (1/(g^2-y^2)) = ?

OpenStudy (jemson):

Hey!!! plz help me.... trying since so many days n recently bumped into this web site... plz let me know the solution

OpenStudy (vishweshshrimali5):

Your first question: (1) \[\large{(x+\cfrac{1}{x})^2 + (x^2 + \cfrac{1}{x^2})^2 + ...+(x^{27} + \cfrac{1}{x^{27}})^2}\] \[\large{= \sum_{i = 1}^{27}(x^i + \cfrac{1}{x^i})^2}\] \[\large{= \sum_{i=1}^{27}((x^i)^2 + 2\times x^i\times \cfrac{1}{x^i} + (\cfrac{1}{x^i})^2)}\] \[\large{= \sum_{i=1}^{27}(x^{2i}) + \sum_{i=1}^{27}2 + \sum_{i=1}^{27}(\cfrac{1}{x^{2i}})}\] \[\large{= S_1 + 2\times 27 + S_2 = S_1 + S_2 + 54}\] Where, \[\large{S_1 = \sum_{i=1}^{27} (x^{2i})}\] \[\large{S_2 = \sum_{i=1}^{27}(\cfrac{1}{x^{2i}})}\] Do you have any doubt till this step ?

OpenStudy (jemson):

i mean its clear.... thnx a bunch

OpenStudy (jemson):

If G is the geometric mean of X and Y, then 1/(G^2 – X^2) + 1/(G^2 – Y^2) =? How to answer this question?.... not getting.

OpenStudy (vishweshshrimali5):

@Jemson , Geometric mean of numbers X and Y , \[\large{G = \sqrt{XY}}\] Thus, \[\large{\cfrac{1}{G^2 - X^2} + \cfrac{1}{G^2 - Y^2}}\] \[\large{= \cfrac{G^2 - Y^2 + G^2 - X^2}{(G^2-X^2)(G^2-Y^2)}}\] \[\large{= \cfrac{2G^2-(X^2+Y^2)}{(G^2-X^2)(G^2-Y^2)}}\] \[\large{= \cfrac{2(XY) - (X^2+Y^2)}{(XY-X^2)(XY-Y^2)}}\] \[\large{= \cfrac{-(X^2+Y^2-2XY)}{X(Y-X)Y(X-Y)}}\] \[\large{= \cfrac{-(X-Y)^2}{-XY(X-Y)^2}}\] \[\large{= \cfrac{1}{XY}}\]

OpenStudy (vishweshshrimali5):

Tell me if you can't understand any step. Okay ?